User Tools

Site Tools


lbaops:lbanov2011:v441apalog

Parkes is using LBADR for this experiment due to Mark5B recording problems encountered in v271j.

Using DAS profile av_16_2_f

Shaun also tried unflipped profile but no fringes seen in either case in RtFC when Jamie S ran At-Pa fringe check.

Suspected possible vex file problem (as initial vex set up for Parkes recording Mark5) so John and Hayley trying correlation on cuppa.. but no luck getting fringes.

JER discovered problem in the morning - wrong synthesiser connected in focus cabin. Fixed at UT 20:41. (This would also have been a fatal problem for previous experiment v271j.)

UT 2137 Windstowed. Back on source 2155.

UT 231840-233200. Recording stopped.

Fringe test to Parkes revealed fringes only in USB. It was determined that LSB channels would be flipped using the current DAS profile.

Shaun changed DAS profile to av16_lu_f.pro at UT 02:19. Should now be correct for this experiment. Should be able to recover earlier data (from 20:41-02:19) at correlator. (Note added at correlation: Not done as flipped channels will cause an error for astrometry.)

Confirmed good fringes in all channels at 03:21 :) :)

UT ~11:39: Windstowed. Briefly came back around 11:53 but then stowed again until end of experiment at 12:00.

lbaops/lbanov2011/v441apalog.txt · Last modified: 2015/12/18 16:38 (external edit)